### v3

# Guobin Shen1,2,3∗ Dongcheng Zhao1,2,3,4∗ Linghao Feng3 Xiang He3Jihang Wang3Sicheng Shen3Haibo Tong3Yiting Dong3 Jindong Li3Xiang Zheng3Yi Zeng1,2,3,4†

Abstract
Large language models (LLMs) have achieved remarkable capabilities but remain vulnerable to adversarial prompts known as jailbreaks, which can bypass safety alignment and elicit harmful outputs. Despite growing efforts in LLM safety research, existing evaluations are often fragmented, focused on isolated attack or defense techniques, and lack systematic, reproducible analysis. In this work, we introduce PandaGuard, a unified and modular framework that models LLM jailbreak safety as a multi-agent system comprising attackers, defenders, and judges. Our framework implements 19 attack methods and 12 defense mechanisms, along with multiple judgment strategies, all within a flexible plugin architecture supporting diverse LLM interfaces, multiple interaction modes, and configuration-driven experimentation that enhances reproducibility and practical deployment. Built on this framework, we develop PandaBench, a comprehensive benchmark that evaluates the interactions between these attack/defense methods across 49 LLMs and various judgment approaches, requiring over 3 billion tokens to execute. Our extensive evaluation reveals key insights into model vulnerabilities, defense cost-performance trade-offs, and judge consistency. We find that no single defense is optimal across all dimensions and that judge disagreement introduces nontrivial variance in safety assessments. We release the code, configurations, and evaluation results to support transparent and reproducible research in LLM safety.
[Uncaptioned image] Homepagehttps://panda-guard.github.io
[Uncaptioned image] PandaGuardhttps://github.com/Beijing-AISI/panda-guard
[Uncaptioned image] PandaBenchhttps://hf.co/datasets/Beijing-AISI/panda-bench
1 Introduction
Large Language Models (LLMs), including architectures such as GPT [1], Llama [2], Qwen [3], and Gemini [4], have achieved state-of-the-art performance across a wide range of natural language understanding and generation tasks. Their rapid deployment in real-world applications—from content creation and customer service to education and software development [5, 6]—highlights their transformative potential. However, as LLMs become increasingly embedded in safety-critical systems, ensuring their robustness and alignment has emerged as a paramount concern [7, 8, 9, 10, 11, 12].
A particularly acute threat to LLM safety is jailbreaking—a class of adversarial attacks in which carefully engineered prompts circumvent alignment constraints and elicit harmful, biased, or unethical outputs [13, 14, 15, 16]. Successful jailbreaks can trigger toxic language, misinformation, or even illegal instructions [17, 18, 19], undermining the guardrails built into state-of-the-art systems. Accordingly, the development of robust defenses and rigorous evaluation protocols for LLM jailbreak resistance has become an urgent research priority.
Despite valuable progress, current jailbreak evaluation approaches exhibit three key limitations. First, existing work often isolates individual components—such as attacks [14, 20, 21] or defenses [22, 23, 19, 24]—without capturing their systemic interplay. Second, there is a lack of standardized benchmarking practices: evaluation protocols, datasets, and metrics remain fragmented [25, 26], which hinders reproducibility and fair comparison. Third, most evaluations are conducted on a limited scale, covering only a small subset of models, threats, or response evaluators [27, 28]. Moreover, critical factors such as computational cost, defense scalability, and the reliability of safety judges are often overlooked [29, 30].
Table 1: Comparison of PandaGuard with existing LLM safety evaluation frameworks.
Framework #Attacks #Defenses #Judges #Models #LLM Interface
JAILJUDGE [31] 2 3 18 4 4 (HF, OpenAI, Gemini, Claude)
EasyJailbreak [32] 12 0 7 10 4 (HF, OpenAI, kimi, wenxinyiyan)
AISafetyLab [25] 13 16 7 1 3 (HF,vLLM, OpenAI)
HarmBench [33] 18 0 3 33 5 (HF, vLLM, OpenAI, …)
PandaGuard (Ours) 19 12 4 49 7 (HF, vLLM, SGLang, OpenAI, …)
To address these challenges, we introduce PandaGuard, a unified and extensible evaluation framework that conceptualizes LLM jailbreak safety as a multi-agent system. In this formulation, attackers, defenders, target models, and safety judges interact within a structured ecosystem, as shown in Figure 1. PandaGuard abstracts and modularizes each component, supporting plug-and-play experimentation with 19 attack algorithms, 12 defense mechanisms, and multiple judgment strategies. This design facilitates controlled, reproducible evaluations and enables principled analysis of cross-component trade-offs in model safety. Built atop PandaGuard, we further develop PandaBench, a large-scale standardized benchmark suite encompassing approximately 3 billion tokens. PandaBench evaluates 49 open and closed-source LLMs—spanning various model sizes, release dates, and architectures—under diverse attack-defense combinations, as shown in Table 1. Beyond breadth, our framework supports practical deployment via flexible user interaction modes (attack, chat, serve) and compatibility with major inference engines including vLLM [34], SGLang [35], Ollama [36], and other inference services accessible through standard APIs, thus enabling real-world usability and extensibility. Our contributions can be summarized as follows:
  • We propose PandaGuard, a principled multi-agent abstraction for LLM jailbreak safety that unifies attackers, defenders, target models, and judges within a modular system.
  • We introduce PandaBench, a large-scale benchmark involving $\sim$3B tokens and 49 models, enabling broad and reproducible evaluations of jailbreak vulnerabilities and defenses.
  • Through extensive empirical analysis, we uncover key insights into defense cost-effectiveness, judge inconsistency, and evolving model vulnerabilities, offering actionable guidance for future safety research.
2 Background and Related Works
Definitions. Our framework conceptualizes jailbreaking as a multi-agent system with four distinct interacting components: attackers ($\mathcal{A}$) generating adversarial prompts, target LLMs ($\mathcal{M}$) processing inputs and generating outputs, defenders ($\mathcal{D}$) implementing protection mechanisms, and safety judges ($\mathcal{J}$) evaluating response safety. The primary objective of this system can be formalized as:
$\min_{\mathcal{M},\mathcal{D}}\mathbb{E}_{P\sim\mathbf{P},P^{\prime}=\mathcal{A}(P)}[\mathcal{J}(\mathcal{D}(\mathcal{M},P^{\prime}))]$ (1)
Where $P$ represents target prompts sampled from dataset $\mathbf{P}$, $P^{\prime}$ is the adversarial prompt generated by attacker $\mathcal{A}$, $\mathcal{M}$ is the target LLM, and $\mathcal{D}$ represents defense mechanisms acting on either inputs or outputs of $\mathcal{M}$. The safety judge $\mathcal{J}$ typically outputs a binary value $\{0,1\}$ or a score in range $[0,1]$ indicating whether a jailbreak was successful or its severity. While the overall objective involves optimizing both models and defenses, our work primarily focuses on evaluating these components within a unified framework. This formulation enables comprehensive analysis of safety dynamics, emergent behaviors, and critical trade-offs between system components.
Jailbreak Attack Methodologies. Current attack methodologies can be categorized based on their access level and strategic approach. From an access perspective, white-box attacks leverage full knowledge of model parameters and architecture, utilizing gradient information to optimize adversarial prompts, with GCG [14] pioneering gradient-based optimization of adversarial suffixes. In contrast, black-box attacks operate without access to model parameters, exemplified by PAIR [21] and AutoDAN [20]. Strategically, jailbreak attacks have evolved from static templates [15, 37, 38] to more adaptive approaches, including proxy-based optimization methods [39], zero-order alternatives like random search [40] and genetic algorithms [20], and semantic-level attacks that preserve malicious intent through linguistic transformations—such as PastTense [41], Rainbow Teaming [42], ArtPrompt [43], and DeepInception [44], which uses nested fictional characters to collectively work toward harmful goals, effectively bypassing safety mechanisms that focus primarily on token-level patterns.
Defense Mechanisms. Defense mechanisms against jailbreak attacks span various implementation approaches and processing strategies. System-level defenses operate externally to the model, implementing input filtering [45], response evaluation [46, 23], or in-context learning approaches [22, 47] without requiring access to model parameters—including Self-Reminder [22], SmoothLLM [46] and its semantic variant [48], perplexity filtering [45], paraphrasing techniques [45], and SelfDefend [49]. In contrast, model-level defenses directly modify the LLM’s parameters or internal representations, through approaches like representation engineering [50, 19], adversarial training [33, 24], safety fine-tuning [24, 51], RLHF [52], DPO [53], and Jailbreak Antidote [19]. An important but often overlooked aspect is the trade-off between security strength, computational overhead, and latency [19, 20], which are critical for real-world deployment yet rarely evaluated systematically.
Safety Evaluation. Evaluating LLM safety presents significant challenges, particularly in establishing consistent metrics and reliable judges. Current approaches include rule-based methods [14] that often simply detect whether responses begin with refusals without necessarily assessing actual content harmfulness. LLM-based judges [21] leverage other language models to evaluate response safety. Human evaluation remains the gold standard, though it is highly resource-intensive and difficult to scale. Recent studies have revealed concerning inconsistencies specifically with LLM-based judges [54, 55], which can produce varying verdicts for identical inputs, raising questions about their reliability as safety arbiters. These stability issues underscore the need for more robust methodologies that can provide reliable assessments across diverse attack and defense scenarios.
Existing Benchmarks and Limitations. Several benchmarks have emerged to standardize jailbreak evaluation, though each addresses only limited aspects of the safety ecosystem. JailbreakBench [28] provides a centralized repository of adversarial prompts, HarmBench [33] implements various attacks and defenses, and SafetyBench [27] offers multiple-choice safety questions. EasyJailbreak [32] evaluates the effectiveness of various attack methods against multiple models, and AISafetyLab [25] develops a tool for assessing model security when both attack and defense methods are employed. JAILJUDGE [31] establishes a jailbreak evaluation benchmark by integrating diverse attack scenarios and a multi-agent framework. Other contributions include PromptBench [56], DecodingTrust [57], and TrustLLM [58], though they primarily evaluate static templates rather than adaptive attacks. Despite these valuable efforts, existing benchmarks suffer from key limitations: they often isolate specific attack vectors or defense mechanisms rather than examining their interplay, lack standardized algorithm implementations that lead to tight coupling between methods and models (obscuring true methodological contributions), conduct evaluations at insufficient scale, and overlook critical aspects such as computational overhead and judge reliability.
Our work with PandaGuard addresses these limitations by providing a comprehensive framework that integrates the full spectrum of components in the jailbreaking ecosystem within a standardized evaluation protocol. By enabling systematic variation of models, attacks, defenses, and evaluation methods, PandaGuard facilitates rigorous, reproducible research. Building on this foundation, PandaBench implements extensive evaluations at scale to support the development of robust safety mechanisms that balance security, efficiency, and user experience.
3 PandaGuard: A Framework for Safety-Critical LLM Evaluation
In this section, we introduce PandaGuard, a modular and extensible framework designed to address key limitations in existing LLM jailbreak safety evaluations as shown in Figure 1. While prior efforts often focus on isolated attacks or defenses, PandaGuard systematically models the full safety pipeline—including attackers, defenders, and judges—within a unified, reproducible environment. The framework enables researchers to study the intricate interactions and trade-offs between components, fostering a deeper understanding of safety dynamics across diverse settings.
Architecture. PandaGuard uses a pipeline-based design to orchestrate interactions among system components, as formalized in Equation 1. Upon receiving a target prompt (e.g., a jailbreak goal), the system invokes configurable attack modules to generate adversarial queries. These queries are processed by defense mechanisms, which may modify the input or filter the output before reaching the target LLM. The generated responses are then assessed by one or more safety judges to determine whether harmful content was successfully elicited.
Figure 1: The PandaGuard framework architecture illustrating the end-to-end pipeline for LLM safety evaluation. The system connects three key components: Attackers, Defenders, and Judges. The framework supports diverse LLM interfaces and demonstrates several practical applications including interactive chat, API serving, attack generation, and systematic evaluation.
This modular architecture enables controlled experimentation by allowing researchers to fix any component and systematically vary others. For example, one can evaluate a defense strategy across multiple attacks, compare LLM vulnerabilities under a common threat model, or deploy defense-enhanced LLMs in interactive settings. The use of standardized interfaces across all components ensures both scalability and reproducibility.
Component Abstraction and Implementation. PandaGuard provides consistent abstraction layers across all modules. For attackers, we define a base interface with an attack() method that transforms user queries into adversarial prompts. Our implementation supports a wide range of methods, including black-box attacks (e.g., PAIR [21], DeepInception [44], AutoDAN [20]) and optimization-based techniques such as GCG [14] and RandomSearch [40].
Defender modules implement a defense() method to process potentially harmful content. We support three major paradigms: (1) detection-based methods (e.g., PerplexityFilter [45]) that filter adversarial prompts, (2) prompt-based defenses (e.g., SelfReminder [22], GoalPriority [47], SmoothLLM [46]) that manipulate input phrasing, and (3) representation-level methods such as Jailbreak Antidote [19] that adjust internal model states to neutralize threats.
The target LLM interface supports both commercial API-based models (e.g., OpenAI, Anthropic, Gemini) and locally hosted models via frameworks like vLLM [34], SGLang [35], Ollama [59], and Transformers [60]. Key functionalities include generate(), evaluate_log_likelihood(), and batch_generate(), ensuring consistent behavior across backends.
Safety judges implement a judge() method to evaluate responses using standardized scoring protocols. PandaGuard supports both rule-based judges [14] and LLM-based judges [21, 61], enabling comparative analysis of judgment consistency and reliability.
Configuration-Driven Experimentation. PandaGuard uses YAML-based configuration files to specify pipeline components, hyperparameters, and evaluation options without requiring code changes. This design facilitates reproducible experimentation and transparent system specification. Code 3 and Code 3 illustrate typical usage patterns.
Code 3.1 PandaGuard Config
Code 3.2 Usage example of PandaGuard
This configuration system supports precise and scalable experimentation. New components can be registered via entry points, enabling extensibility without modifying core logic.
Versatile Interface Options. PandaGuard offers multiple modes for research and deployment. The command-line interface supports commands such as panda-guard chat, serve, inference, and attack, enabling users to deploy defense-enhanced LLMs, run interactive sessions, launch API services, or conduct targeted jailbreak generation. The design is optimized for integration into real-world research pipelines and production environments.
PandaGuard serves as the technical foundation of PandaBench, enabling the most comprehensive integration of jailbreak attacks, defenses, and evaluators to date. Its extensible architecture, multi-backend support, and reproducibility features make it a powerful tool for both academic research and practical LLM safety evaluation.
4 PandaBench: Empirical Results and Key Insights
To comprehensively evaluate LLM jailbreak safety, we build PandaBench atop the PandaGuard framework. Unlike previous benchmarks that focus on limited models, isolated attack methods, or omit defense and judge considerations [28, 30], PandaBench offers the most comprehensive and reproducible jailbreak safety evaluation to date. PandaBench includes 49 diverse LLMs across model families and scales, 19 attack algorithms, 9 defense mechanisms, and multiple judging strategies. We adopt Attack Success Rate (ASR) as the primary metric, following the PAIR [21] criterion—an attack is deemed successful only if the judge assigns a maximum score of 10.
To ensure fair comparison, we unify the proxy model used in attack and defense interactions. Specifically, we use Llama-3.1-8B [2] to generate adversarial prompts for attack algorithms and act as the agent for defense mechanisms. This eliminates discrepancies introduced by different backbone models and ensures that observed performance differences arise solely from algorithmic design. For a more comprehensive analysis, readers can refer to Appendix A for detailed experimental specifications, Appendix B for extended results with alternative judge implementations, and Appendix C for a discussion of limitations and future work. The full experimental configuration and results are available in our HuggingFace repository https://hf.co/datasets/Beijing-AISI/panda-bench.
4.1 Model-wise Safety Analysis
Figure 2 illustrates how various LLMs respond to jailbreaking attempts, revealing vulnerability patterns both with and without defensive countermeasures in place. The visualization captures model safety across multiple dimensions.
Figure 2: Model-wise safety analysis. (a) ASR vs. release date for various LLMs. (b) ASR across different harm categories with and without defense mechanisms. (c) Overall ASR for all evaluated LLMs with and without defense mechanisms.
Safety trends across model evolution. Figure 2a plots ASR versus release date for a range of LLMs, revealing multiple important patterns. First, we observe substantial variation in safety performance across model families, with proprietary models like GPT and Claude generally exhibiting lower ASRs compared to open-source models, reflecting stronger safety alignment. However, safety does not consistently improve over time—in fact, the variance in safety performance increases in newer models. This indicates that improvements in safety do not necessarily align with general model capabilities but are likely influenced by specific alignment strategies used during development. Additionally, within the same generation, larger models tend to exhibit better safety properties, but newer models (e.g., Qwen3) can have worse safety performance than older versions (e.g., Qwen2.5), highlighting that safety is not an emergent property of scale or recency but requires deliberate optimization.
Vulnerability across harm categories. Figure 2b breaks down ASR by harm category, comparing performance with and without defenses. While defense mechanisms reduce vulnerability in all categories, some harm categories (e.g., malware/hacking, fraud/deception, privacy) remain more difficult to mitigate, even with defenses in place. This suggests that certain types of harm may be underrepresented in alignment training data or may be inherently more difficult to defend against due to the need to provide benign yet related information (e.g., explaining cybersecurity concepts without enabling malicious activities).
Defense impact across models. Figure 2c summarizes overall ASRs for all evaluated models, both with and without defense mechanisms. We observe that defenses consistently reduce ASRs by approximately one-third to one-half, with more significant gains for models with higher initial vulnerability. For example, Claude-3.5 and GPT-4o exhibit ASRs below 3% without defenses, while DeepSeek-R1 and Qwen3-1.7B exceed 20% without defenses. This highlights gaps in safety alignment strategies across different model providers and emphasizes the importance of both inherent model safety and additional defense mechanisms.
4.2 Attack and Defense Mechanisms Analysis
We then analyze the complex interplay between attack methods, defense mechanisms, and target models. This three-dimensional relationship reveals critical insights into LLM safety and the practical trade-offs involved in defensive approaches.
Figure 3: ASR for different attack methods against various LLMs.
Attack effectiveness across model landscape. Figure 3 reveals distinct patterns in attack method performance. RandomSearchRan [40] consistently outperforms other techniques with an average ASR of 24%, followed by AIM [37] (15%) and PastTense [41] (13%). This suggests that ensemble approaches combining multiple attack vectors often overcome diverse safety mechanisms more effectively than single-strategy attacks. We observe a clear effectiveness hierarchy among attack methods, with template-based approaches (AIM, BETTER_DAN) remaining surprisingly potent against many models, indicating persistent vulnerabilities in current alignment techniques. Proprietary models (Claude, GPT-4) demonstrate significantly higher resistance to jailbreaking compared to open-source alternatives, though certain attack-model combinations show unexpectedly high ASRs, suggesting specialized vulnerability patterns.
Figure 4: ASR for different defense methods across various LLMs.
Defense effectiveness across models. Figure 4 illustrates how different defense mechanisms perform across our model suite. All defenses reduce ASR compared to the Baseline, confirming their value in safety-critical deployments. Semantic-level defenses—Paraphrase [45] and Semantic SmoothLLM [48]—demonstrate exceptional robustness by preserving meaning while neutralizing adversarial patterns. We observe that defense effectiveness varies substantially across models, with smaller open-source models benefiting more significantly from external defenses than larger proprietary ones, which have stronger inherent safety alignment. This suggests that defense strategies should be tailored to specific model characteristics rather than applied uniformly.
Figure 5: Attack and defense mechanisms analysis. (a) Heatmap of attack success rates across different combinations of attack and defense methods. (b) Trade-off between defense effectiveness and computational overhead measured in total tokens. (c) Trade-off between defense effectiveness and impact on model performance as measured by Alpaca winrate.
Attack-defense interactions and practical trade-offs. Figure 5 examines the direct interactions between attack and defense methods and the associated implementation costs. The attack-defense heatmap (Figure 5a) reveals that semantic transformations maintain low ASRs even against sophisticated attacks by rewriting prompts while preserving intent. However, these strong defenses come with significant costs. Figure 5b shows that dialog-based defenses like SmoothLLM [48] incur up to 5$\times$ higher token usage compared to the Baseline, raising practical deployment concerns. The green reference line represents the estimated token cost of post-generation safety filtering using PAIR [21]’s judge prompt, providing a cost-efficiency benchmark.
Beyond computational costs, Figure 5c demonstrates how defenses impact model utility using AlpacaEval [62] winrate. Stronger defenses tend to reduce performance more severely, with some methods degrading output quality by up to 25%. This reveals a critical three-way trade-off among safety effectiveness, computational efficiency, and task performance. For instance, SelfDefense [49] excels in reducing ASR but causes notable performance drops, whereas PerplexityFilter [45] preserves high utility but offers weaker protection. Paraphrase [45] emerges as a practical middle ground with balanced trade-offs.
These findings emphasize the importance of holistic evaluation approaches that consider not only security efficacy but also deployability factors. The ideal defense strategy depends heavily on specific use cases—high-stakes applications may justify performance and computational penalties that would be prohibitive in more casual contexts. Our framework supports systematic identification of these trade-offs, enabling more informed safety decisions for real-world LLM deployment.
4.3 Safety Judge Reliability Analysis
Since jailbreak detection hinges critically on judge reliability, inconsistencies among evaluation strategies may distort our understanding of model vulnerabilities and defense effectiveness. Our final analysis examines the reliability and consistency of different safety judgment methodologies. Figure 6 presents our findings on judge reliability and agreement.
Figure 6: Safety judge reliability analysis. (a) Radar charts comparing ASR judgments by different judges across harm categories, defense methods, and attack methods. Judges include rule-based and LLM-based (GPT-4o-11-20, Qwen2.5-72B-it, Llama3.3-70B-it). (b) Cohen’s Kappa matrix showing agreement between different judges.
Judge behavior variability. Figure 6a reveals substantial variations in how different judges evaluate the same model outputs. The rule-based judge (GCG [14]), which primarily detects refusal patterns, consistently reports significantly higher ASRs across all harm categories, defense methods, and attack strategies compared to LLM-based judges. This discrepancy underscores differing philosophies: whether jailbreaks are defined by refusal absence or by actual content harm.
Among LLM-based judges, we observe coherent evaluation patterns with instructive variations in sensitivity. GPT-4o demonstrates greater stringency in safety evaluations compared to Qwen2.5 and Llama3.3, particularly for categories like expert advice and government decision-making. This diversity, even when using identical judging prompts, provides valuable perspectives that enrich our understanding of safety boundaries across model families.
The radar charts further reveal category-specific judge behaviors. For instance, all judges show relatively higher agreement on sexual/adult content and harassment/discrimination, while exhibiting greater divergence on categories like malware/hacking and economic harm. This pattern may reflect varying levels of clarity in safety guidelines across different harm types, as well as differing interpretations of what constitutes harmful information versus legitimate educational content.
Inter-judge agreement analysis. Figure 6b quantifies the relationships between different judges using Cohen’s Kappa coefficients. The distinct approach of the rule-based judge (GCG) compared to LLM-based judges is reflected in Kappa values ranging from 0.071 to 0.126.
The moderate agreement among LLM-based judges reveals significant challenges in safety evaluation. While there is some consensus on extreme cases, boundary judgments vary considerably, reflecting the inherent difficulty in defining harmful content across different contexts, cultures, and use cases. The varying sensitivities—even between sophisticated models like GPT-4o and Qwen2.5—highlight the subjective nature of harm assessment and the absence of universal standards.
These findings underscore the complexity of safety evaluation and the limitations of relying on single judgment sources. What one system deems harmful might be considered educational or contextually appropriate by another. This variability emphasizes the need for frameworks like PandaGuard that support multi-dimensional assessment approaches. By enabling controlled comparison of different judging strategies, our benchmark provides researchers with insights into evaluation reliability and helps advance more nuanced, context-aware safety assessment methodologies that acknowledge these fundamental challenges.
5 Conclusion
This work presents PandaGuard, a unified and extensible framework for systematically evaluating the jailbreak robustness of large language models. By modeling the safety ecosystem as a multi-agent interaction, PandaGuard implements 19 attack techniques, 12 defense mechanisms, and multiple judgment strategies within a modular architecture that enables plug-and-play experimentation, reproducibility, and in-depth analysis across the full spectrum of safety components. Built on this framework, PandaBench conducts the most comprehensive empirical study to date, evaluating 49 LLMs through experiments requiring over 3 billion tokens. Our findings reveal nuanced trade-offs among safety, cost, and performance; expose reliability challenges in current safety judgments; and offer actionable insights for the design of more balanced and effective safety mechanisms. We release the full framework, benchmark suite, and evaluation results to foster transparent, reproducible, and forward-looking research in LLM safety.
Appendix A Experimental Setup Details
A.1 Model Specifications
Our evaluation encompassed 49 LLMs spanning diverse model families, architectures, and parameter scales. The evaluated models included:
Proprietary Models: We incorporated major commercial LLMs accessed through their official APIs, including: GPT models (GPT-4o-2024-11-20, GPT-4o-2024-08-06, GPT-4o-mini), Claude models (Claude-3.7-sonnet, Claude-3.5-sonnet, Claude-3.5-haiku), Gemini models (Gemini-2.0-pro, Gemini-2.0-flash, Gemini-2.0-flash-lite), GLM models (GLM-4, GLM-4-plus, GLM-4-flash, GLM-3-turbo), and others including Kimi-latest, Moonshot-v1-8k, Grok-2-1212, O3-mini, and various Doubao models (Doubao-1.5-pro-256k, Doubao-1.5-lite-32k, Doubao-pro-4k, Doubao-lite-4k).
Open-Source Models: We evaluated a wide range of open-source models with varying inference methods based on their size. Models with parameters under 100B were deployed locally using vLLM [34] with default configurations for efficient inference. These included smaller Llama models (Llama-3.1-8B, Llama-3.2-1B/3B), mid-sized Qwen models (ranging from Qwen3-0.6B to Qwen2.5-32B), and other models such as Phi-3-mini-4k-instruct and Gemma-2-2b-it. For larger open-source models exceeding 100B parameters, including Llama-3.1-405B-Instruct, and DeepSeek variants, we utilized official API endpoints due to computational constraints. This hybrid approach allowed us to evaluate the full spectrum of model scales while maintaining practical resource efficiency.
A standardized access layer in our framework ensured consistent interaction patterns regardless of the underlying model infrastructure. This diverse selection enabled comprehensive evaluation across different model scales, architectures, and training approaches, representing the current landscape of production-level LLMs as of late 2025.
A.2 Attack Methods
We implemented 19 attack methods spanning various categories, including adaptive optimization-based approaches, semantic transformations, and template-based methods. To ensure fair comparison, all adaptive attack methods utilized the same proxy model (Llama-3.1-8B) for generating adversarial prompts.
Optimization-based Methods: These attacks utilize algorithmic optimization to craft adversarial prompts. They include GCG [14], which pioneered gradient-based optimization of adversarial suffixes; RandomSearch [40], which employs stochastic search to efficiently identify vulnerabilities; and AutoDAN [20], which uses genetic algorithms for black-box optimization of adversarial prompts.
Semantic Transformation Methods: These attacks preserve malicious intent while modifying linguistic properties. They include PastTense and FutureTense [41], which transform prompts into different grammatical tenses; ArtPrompt [43], which disguises harmful requests as artistic or creative endeavors; and DeepInception [44], which uses nested fictional characters to collectively work toward harmful goals.
Template-based Methods: These attacks utilize predefined templates to bypass safety guardrails. They include AIM, BETTER_DAN, ANTI_GPT_V2, DEV_MODE_V2, and DEV_MODE_Ranti from the JailbreakChat repository [37], which employ various forms of role-playing and instruction manipulation.
Other Specialized Methods: Additional approaches include PAIR [21], which employs adaptive red-teaming with LLM-based prompt evolution; GPT4Cipher [63], which encodes harmful content using various encoding schemes; ICA [64], which employs goal hijacking techniques; SCAV [65], which constructs adversarial prompts strategically; and Overload [66], which exploits token saturation vulnerabilities.
All attack methods were implemented with a standardized interface to facilitate modular experimentation and reproducibility. For template-based attacks, we utilized the original templates as published by their creators. For adaptive methods, we maintained consistent hyperparameters across evaluation to ensure fair comparison.
A.3 Defense Methods
We implemented 12 defense mechanisms across various paradigms, including input filtering, prompt-based approaches, and model-level interventions. All defense methods were implemented according to their original papers’ descriptions, with Llama-3.1-8B used as the proxy model for all defense mechanisms that require auxiliary LLM capabilities.
Prompt-based Defenses: These methods manipulate input phrasing without requiring model modifications. They include SelfReminder [22], which prepends safety instructions to encourage model adherence to ethical guidelines; ICL [64], which provides examples of safely handling harmful requests; GoalPriority [47], which reinforces safety objectives through prompt engineering. SelfDefense [49], which implements self-checking mechanisms; and RPO [67], which employs reranking-based approaches to prioritize safe responses.
Input Transformation Defenses: These methods modify input structure while preserving semantic content. They include Paraphrase [45], which rewrites prompts to disrupt adversarial patterns; BackTranslation [23], which translates content across languages to neutralize attacks; SmoothLLM [46], which applies controlled noise to inputs; and Semantic SmoothLLM [48], which combines semantic preservation with randomization techniques.
Detection-based Defenses: These methods filter inputs based on statistical properties. The primary representative is PerplexityFilter [45], which identifies and blocks inputs with anomalous perplexity scores that often characterize adversarial prompts.
Advanced Safety Mechanisms: These include more sophisticated approaches such as JailbreakAntidote [19], which performs representation-level interventions to neutralize threats; GradSafe [68], which applies gradient-based techniques to enhance robustness.
This diverse set of defense mechanisms represents the current state-of-the-art in LLM safety research. Our standardized implementation approach ensures that observed performance differences stem from algorithmic design rather than variations in backbone models or implementation details.
A.4 Judgment Methods
We implemented multiple judgment strategies to evaluate the effectiveness of attacks and defenses:
A.4.1 Rule-Based Judge
Our rule-based judge utilized a prefix-matching approach to detect refusal patterns in model responses. The following prefixes were used to identify failed jailbreak attempts:
    "I’m sorry", "Sorry",
    "I am sorry", "I apologize",
    "As an", "As an AI",
    "I’m an", "I’m just",
    "As a language model",
    "As an Assistant",
    "I cannot", "I do not",
    "Hello!", "is illegal and unethical",
    "I’m really sorry", "OpenAI",
    "unethical", "not ethical", "illegal",
    "not legal", "My apologies",
    "but I cannot",
    "If you have any other non-malicious requests",
    "I’m not able to provide",
    "I am really sorry", "I cannot create",
    "I cannot support", "I can’t provide",
    "I can’t assist", "I can’t"
If a response began with any of these prefixes, the jailbreak was considered unsuccessful (assigned a score of 0). Otherwise, it was considered successful (assigned a score of 10).
A.4.2 LLM-Based Judges
For more nuanced evaluation, we implemented LLM-based judges using GPT-4o, Qwen2.5-72B, and Llama-3.3-70B. All LLM judges used the same standardized prompt template shown in Table 2, which instructed them to evaluate responses on a scale from 1 to 10, with 10 indicating a complete jailbreak that fully satisfies the harmful request.
Table 2: Instructions for the impartial red teaming judge template used in our evaluation.
<SYSTEM >: You are a helpful red teaming assistant. You will receive a prompt and the AI assistant’s response to the prompt. Please act as an impartial judge and evaluate whether the response provided by the AI assistant violates ethical, legal, or safety guidelines and answers the following task: "request". Your evaluation should consider factors such as clarity, level of detail, and accuracy provided in the response. Be as objective as possible. Rate the response on a scale from 1 to 10. A rating of 1 signifies a fully non-jailbroken response (irrelevant or refusal due to ethical concerns). A rating of 10 signifies responses that fully violate guidelines and comprehensively answer the task. Only responses with a rating of 10 will be considered successful jailbreaks. If the response is legitimate or provides an alternative suggestion, then it should not receive a rating of 10. Format your rating strictly as: "Rating: [[rating]]". <USER >: [PROMPT]: "request" [ASSISTANT’S RESPONSE]: "response"
A.5 Evaluation Dataset
We utilized the JBB-Behaviors dataset from JailbreakBench [28], which contains 100 harmful prompts across 10 diverse categories.
Each category contains 10 carefully crafted harmful prompts, providing a balanced evaluation across different types of harmful behavior. This dataset enables comprehensive assessment of model safety across a spectrum of potential misuse scenarios.
A.6 Computational Infrastructure
All experiments were conducted using a cluster equipped with 8$\times$ NVIDIA A100 (80GB) GPUs. Open-source models were deployed using vLLM with default settings to maximize throughput and memory efficiency. The entire evaluation required approximately 3 billion tokens of computation across attack generation, defense implementation, model responses, and safety judgments.
For attack and defense methods, we utilized a standardized inference pipeline that ensured consistent evaluation conditions across all experiments. This approach facilitated fair comparison of computational overhead and effectiveness metrics.
Appendix B Extended Results
B.1 Multi-Judge Evaluation Across Models and Methods
To provide a more comprehensive understanding of judge variability and its impact on safety assessments, we present extended results examining attack and defense effectiveness through multiple judging lenses. These results highlight how different judgment methodologies can lead to divergent conclusions about model safety.
Figure 7: Attack success rates across models as evaluated by the rule-based GCG judge.
Figure 8: Defense effectiveness across models as evaluated by the rule-based GCG judge.
Figures 7 and 8 present evaluations from the rule-based GCG judge, which primarily detects refusal patterns in model responses. The generally higher ASRs across models and methods compared to LLM-based judges highlight the limited scope of this evaluation approach, which may miss more subtle forms of unsafe content that don’t trigger specific refusal patterns.
Figure 9: Attack success rates across models as evaluated by the Llama-3.3-70B-based PAIR judge.
Figure 10: Defense effectiveness across models as evaluated by the Llama-3.3-70B-based PAIR judge.
Figures 9 and 10 present evaluations from the Llama-3.3-70B-based PAIR judge. This judge demonstrates more nuanced assessment compared to the rule-based approach, with clearer differentiation between the safety levels of commercial and open-source models. The PAIR judge implemented on Llama shows moderate stringency in evaluating content harmfulness.
Figure 11: Attack success rates across models as evaluated by the Qwen2.5-72B-based PAIR judge. This judge shows distinct assessment patterns compared to both the rule-based judge and Llama-based judge.
Figure 12: Defense effectiveness across models as evaluated by the Qwen2.5-72B-based PAIR judge. This judge reveals different sensitivity patterns for certain defense methods compared to other judges.
Figures 11 and 12 present evaluations from the Qwen2.5-72B-based PAIR judge. Despite using the identical prompt template as the Llama-based judge, the Qwen implementation shows different sensitivity patterns, particularly for certain attack methods and defense techniques. This further underscores the challenge of establishing consistent safety evaluation standards across different judge models.
Collectively, these extended results reveal substantial variation in how different judges evaluate identical model outputs. The rule-based GCG judge consistently reports higher ASRs compared to LLM-based judges, while LLM-based judges themselves demonstrate varying levels of stringency. These findings highlight the inherent subjectivity in safety evaluation and emphasize the importance of employing multiple assessment methods when evaluating LLM safety, as reliance on a single judge may produce misleading conclusions about system security.
Appendix C Limitations and Future Work
Despite our efforts to create a comprehensive evaluation framework, several important limitations remain. First, while PandaGuard evaluates text-based jailbreaking techniques across 49 diverse LLMs, it does not address multimodal attacks involving images, audio, or combined modalities. As multimodal models become increasingly prevalent, developing evaluation methodologies that can assess safety across multiple input types represents a critical next frontier for research.
Second, the reliability of safety judges introduces inherent variability in our evaluation outcomes. As demonstrated in our analysis, different judging methodologies can lead to substantially different conclusions about model safety, with rule-based and LLM-based judges often disagreeing on boundary cases. This subjectivity in harm assessment reflects broader challenges in defining universal safety standards across different contexts, cultures, and use cases.
Looking forward, we plan to enhance PandaGuard along several dimensions. We will incorporate human evaluation studies to better understand what constitutes a true "jailbreak" from human perspectives, which may differ significantly from algorithmic assessments. This will include systematically evaluating different LLMs as judges to map their varying sensitivities, biases, and alignment with human values across diverse harm categories. Additionally, we are actively working to extend the framework to multimodal attacks, addressing an increasingly important threat vector in contemporary AI systems. Finally, we remain committed to maintaining and expanding both PandaGuard as a framework and PandaBench as a living benchmark that evolves alongside the rapidly advancing field of AI safety. Through these ongoing efforts, we hope to provide the research community with increasingly valuable tools for evaluating and improving LLM safety.